OpenStudy (anonymous):

Find the unit rate. Round to the nearest tenth, if necessary. 344 miles in 11 hours

OpenStudy (anonymous):

\[\frac{344 [miles]}{11 [hours]}=\frac{344}{11}\frac{miles}{hour}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

what did you end up rounding to?